Fermi National Laboratory


Accelerator Update for September 30, 2005 - October 3, 2005

Friday September 30
The day shift began with Operations monitoring a Recycler stash, with the Tevatron (TeV) off for access, with the Antiproton Source (Pbar) stacking, with the Collider Detector at Fermilab (CDF) on access to their collision hall, with D-Zero (D0) waiting for the next TeV store, and with the MiniBooNE, NuMI, and the Meson MCenter experiments taking beam.

CDF came out of access at 8:05 AM.

Operations put the TeV through a dry squeeze at 8:43 AM.

At 9:32 AM, Main Injector RF stations (MIRF4 & 6) tripped off. They reset okay.

At 9:32 AM, Operations reported that NuMI tripped off due to a number of power supplies tripping off coincidentally with MIRF4 & 6 tripping off. Everything reset okay.

An operator shut off beam to MiniBooNE at 9:38 AM so that experts could access MI-13. They came out of access and had beam reestablished about an hour later.

Operations put the TeV into shot setup at 10:18 AM.

At 11:56 AM, Operations reported that the Recycler stash disappeared due to a few corrector magnet bulk power supplies tripping off on an over temperature indication. An operator informed an expert.

At 12:37 PM, the Cryo System Coordinator reported that a TeV sector A4 cryo valve (A4EVLH) was stuck. Operations had already loaded all the protons and most of the antiprotons into the TeV for the next store. The setup was put on hold.

Cryo system technicians reported at 12:41 PM that they had used another valve to control the system, which allowed Operations to continue with shot setup. Later, the cryo techs replaced the actuator for EVLH.

Operations established store 4414 at 1:15 PM with an initial luminosity of 66.58E30. Pbar resumed stacking and MCenter resumed taking beam.

At 2:05 PM, techs at the Central Utility Building (CUB) reported a power glitch that knocked off their chillers. Operators investigated but couldn't find any other indications of a power glitch.

An expert turned on the Recycler bulk power supplies at 4:26 PM.

Operations began preparing the Recycler to stash at 8:32 PM. The stash ended and Pbar and MCenter resumed normal operations.

Saturday October 1
The midnight shift began with Operations monitoring store 4414 and the stash, with Pbar stacking, with NuMI, MiniBooNE, and MCenter taking beam.

At 1:14 AM, Operations reported that the NuMI horn tripped off on a ground fault indication. It would reset, but then trip off again. Operations called in experts.

Operations terminated store 4414 at 10:18 AM.

Operations prepared D0 for a collision hall access at 10:48 AM.

At 11:12 AM, MIRF11 tripped off multiple times over the next twenty minutes. Operations contacted an expert who came in and made adjustments.

D0 came out of access at 11:33 AM.

Operations put the TeV through a dry squeeze at 12:03 PM.

Operations put the TeV into shot setup at 12:45 PM.

At 3:49 PM, just after all the protons and antiprotons had been loaded and squeezed, the TeV suffered a sector A1 quench.

NuMI experts reported around 7 PM that they had a dead short in the tunnel. They reported that they would access the tunnel in the morning and continue their investigations.

TeV experts reported at 7:36 PM that they had found a vertical distortion near B0. They suggested a procedure for Operations to follow before the next shot.

Operations put the TeV through a dry squeeze at 8:31 PM.

Operations put the TeV through a wet squeeze at 9:15 PM.

Operations began an orbit smooth at around 11:30 PM.

Sunday October 2
The midnight shift began with Operations preparing the TeV for the next store, with Operations monitoring the stash, with Pbar stacking, with NuMI off due to a possible Horn ground fault, and with MiniBooNE and MCenter taking beam.

Operations put the TeV into shot setup at 2:24 AM.

At 4:22 AM, the TeV suffered a sector D1 quench.

Pbar resumed stacking at 4:31 AM.

MCenter resumed taking beam at 4:54 AM.

Operations reported at 6:18 AM that the TeV sector B17 separator hadn't put out the correct voltage and that had caused the quench. An operator called in a power supply expert.

Operations put the TeV through a dry squeeze at 8:14 AM.

A TeV expert reported at 9:07 AM that the B17 separator appeared to have a bad cable in the tunnel.

Operations prepared the TeV for access at 9:24 AM. Besides investigating the B17 separator, two other TeV teams worked on a vacuum leak at F0 and on vacuum pumps at A14 and B43.

Operations began preparing the Recycler at 11:12 AM to stash antiprotons. The stash was completed by 11:50 AM and Pbar resumed stacking.

MCenter resumed taking beam at 12:16 PM.

The TeV access ended at 12:55 PM.

At 2:49 PM, Operations reported that a lightning strike took down the Booster RF, the MI power supply loop, and the TeV RF monitor.

A NuMI expert reported at 2:52 PM that they verified the horn fault. They began isolating Horn #2.

Operations put the TeV through a dry squeeze at 4:31 PM.

Operations loaded a 12x0 study-store into the TeV at 6:07 PM. They conducted a wet squeeze at 6:22 PM.

Operations reported at 6:53 PM that they had completed the orbit corrections.

Somewhere around 8 PM, Operations put the TeV into shot setup.

Operations established combination store 4428 at 10:34 PM with an initial luminosity of 106.44E30. Pbar resumed stacking.

Monday October 3
The midnight shift began with Operations monitoring store 4428, with Pbar stacking, with NuMI off due to a Horn ground fault, and with MiniBooNE and MCenter taking beam.

The Recycler began preparing to stash antiprotons at 3:43 AM. The stash ended and MCenter resumed taking beam at 4:31 AM.

Pbar resumed stacking at 4:38 AM.

The Plans for Monday
The plans for today are to stack, stash, and store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 10/3/2005   email Fermilab